OpenStudy (anonymous):

1. One of the solutions to the quadratic equation x2 - 13x + 30 is 3 whats another? 2. and one of the solutions to the quadratic equation 2x2 - 9x – 5 is -1/2 whats also another ?

OpenStudy (anonymous):

Are you just checking your answer, or do you want an explanation?

OpenStudy (anonymous):

well im just checking and i need to write down two for each so im unsure whats the other

OpenStudy (campbell_st):

factorise the quadratic factors of 30 that add to -13....... -10, -3 (x -3)(x-10) = 0 you can finish

OpenStudy (anonymous):

first one is 10, second one is 5.

OpenStudy (campbell_st):

(2) 2 x -5 = -10 find the factors of -10 that add to -9.... -10 and 1 (2x -10)(2x+1)/2 = 2(x - 5)(2x + 1)/2 gives (x - 5)(2x + 1) = 0 you can finish this one

OpenStudy (anonymous):

The product of the two roots of \[ a x^2 + b x + c=0\] is \[\frac c a\] for the first quadratic the product of the two roots is \[\frac c a = \frac {30}1 = 30 \] If one of the roots is 3, what is the other one.

OpenStudy (anonymous):

That's interesting, I didn't know there was a formula for the product of the roots.

OpenStudy (anonymous):

For the second one \[ \frac c a = - \frac 5 2 \] The first root is \[ -\frac 1 2 \] The second must be 5.

OpenStudy (campbell_st):

sum of the roots \[\alpha + \beta = -b/a\]

OpenStudy (anonymous):

@campbell_st, you can do it this way two. Problems like this can be found on http://www.saab.org/mathdrills/act.html